# Non-Abelian anyons in a periodically-driven Abelian model

### 06 Feb 2024 v1

@@ Abstract
Floquet-Majorana modes. For weak driving, the fermions reproduce faithfully the behaviour of a topological superconductor, as described by the time-averaged Hamiltonian (5). We study these properties by numerically computing the exact Floquet Bogoliubov-de Gennes (FBdG) Hamiltonian $\hat{H}_{\rm FBdG}$, defined by $\hat{H}_{F}=(1/2)(\hat{\bm{f}}^{\dagger},\hat{\bm{f}})\hat{H}_{\rm FBdG}(\hat{\bm{f}},\hat{\bm{f}}^{\dagger})^{T}$, as well as its quasienergy spectrum and ground state parity, from the driven Hamiltonian of Eq. (3[22]. This is done through the diagonalization of the one-cycle evolution operator $\hat{U}(T)$ and of the quasienergy operator in Floquet space [30]. In the absence of vortices [$\mathcal{S}_{p}^{(r)}=1$], we find four nearly-degenerate Floquet high-frequency ‘ground’ states, one in each superselection sector corresponding to periodic or anti-periodic boundary conditions for the fermions in each direction [15], as given by the absence or presence of a $e$ particle in the holes of the torus. The topological phase, characterized by a non-zero Chern number of the negative-energy fermion band, is expected for $|\mu|<4J$ [15, 32, 24]. In such a phase, only three ground states have even fermion parity and are thus physical, corresponding to the three Ising-anyon types $e,1,\psi$. Choosing the driving frequency $\omega$ such that the effective chemical potential is in the topological region, $\mu\simeq-2J$, the fourfold degeneracy is approached exponentially as the system size increases, as shown in Fig. (2)(a), and agrees wih the predicted fermion parities. On a cylinder, the quasienergy spectrum exhibits two edge modes [Fig. 1(b)], signalling the achievement of non-zero Chern number, which disappear for $|\mu|>4J$.
@@
When introducing two vortices at separate positions on the torus, approximate Majorana modes at near-zero quasienergy appear: they persist as $\omega$ is detuned from resonance within the topological phase and are exponentially localized around the vortices [Fig. 2(d)]. This confirms that the high-frequency description of Eq. (5) correctly captures the topological properties of the system, which are not spoiled by the higher-order contributions to the high-frequency expansion neglected in (5). Those terms, besides renormalizing the couplings in $\hat{H}_{F}$, generically contain coupling terms of progressively longer range [30, 33]. Increasing the driving amplitudes $J=\Delta$, while keeping $\mu=-2J$ to stay in the topological phase, the driven fermions enter an anomalous Floquet-topological phase [34, 35, 26] as they depart from the high-frequency limit. Although, in this regime, the description in terms of the time-averaged Hamiltonian of Eq. (5) is no longer valid, the system still exhibits Majorana modes, at both quasienergy zero and ±$\pm\pi/T$, see Fig. 2(c). This effect is only possible in driven systems [34, 36, 37, 38] and, in the context of Kitaev quantum wires, has been proposed as a means to realize non-Abelian braiding and quantum computation with edge modes in 1D [39, 40, 41, 42].
Figure 3: (a) Paths $P$ and $P^{\prime}$ involved in the vortex-exchange protocol. The phase difference between the two cases determines the exchange phase. (b) Vortex exchange phase $\theta_{P}-\theta_{P^{\prime}}$ for $\omega=2(\mu_{\psi}-2J)$, different system size and driving amplitude in the high-frequency regime. The upper (lower) panel shows the phase for doubly-periodic (anti-periodic) fermion boundary conditions, corresponding to the $\psi$ ($\mathds{1}_{\rm is}$) sector.
@@
Non-Abelian exchange phases. We next analyze the exchange statistics of $e$ vortices in the fermion topological phase, to verify whether they realize non-Abelian statistics as predicted in the presence of Majorana modes [8, 23]. In the weak-driving regime, the ground state of the time-averaged effective Hamiltonian may be prepared adiabatically from the ground state of the undriven model by slowly ramping up the Floquet modulation [43, 44], which also averages out micromotion effects related to the global phase of the drives [29]. We exchange a pair of vortices using the approach of Ref. [21], based on Levin-Wen’s protocol [45, 8, 46] and sketched in Fig. 3(a). Such a protocol compares the phase accumulated by moving the vortices along two paths $P$ and $P^{\prime}$ sharing the same set of vortex positions, but where the vortices are moved in different order. This guarantees that the Berry-phase difference between the two paths is determined by the exchange statistics only. With the allowed control on the toric-code spins, the operators moving a vortex between nearby vertices $v$ and $v+r$ are the Pauli operators $\hat{Z}_{v,R}$ (for $r=x$) and $\hat{Z}_{v,B}$ (for $r=y$) acting on the spin shared between them. In the quasiparticle picture, the latter map to quasiparticle operators $\hat{t}_{v}^{(r)}=\hat{\tilde{\mathcal{S}}}_{v}^{(r)}(\hat{b}_{v}^{\dagger}+\hat{b}_{v})(\hat{b}_{v+r}^{\dagger}+\hat{b}_{v+r})$, inducing boson motion and also containing a fermionic part to account for semionic statistics, $\hat{\tilde{\mathcal{S}}}_{v}^{(x)}=1$ and $\hat{\tilde{\mathcal{S}}}_{v}^{(y)}=\prod_{p\in\mathrm{L}[p(v+y)]}(-1)^{\hat{f}_{p}^{\dagger}\hat{f}_{p}}$, where $\mathrm{L}[p(v+y)]$ denotes all plaquettes to the left of the spin shared by $v$ and $v+y$. A vortex may be moved by locally realizing the single-spin transformation $\hat{Z}_{v,\alpha}$ through a Floquet-adiabatic process. For each intermediate vortex configuration |$\ket{\vec{n}_{i}^{e}}$ along the exchange paths, the fermions realize the corresponding BCS even-parity ground state |$\ket{\Phi_{i}}$ of the FBdG Hamiltonian, which we determine numerically. The Berry phase accumulated along the path $P$ is computed as $\theta_{P}=\mathrm{Arg}\prod_{(i,i+1)\in P}\bra{\Phi_{i+1}}\hat{\tilde{S}}^{(r)}_{i+1,i}\ket{\Phi_{i}}$, where we indicated with $\hat{\tilde{S}}^{(r)}_{i+1,i}$ the fermionic part of the vortex-displacement operator $\hat{t}_{v}^{(r)}$ which converts the vortex configuration |$\ket{\vec{n}^{e}_{i}}$ into |$\ket{\vec{n}^{e}_{i+1}}$ by moving one vortex. Details about the numerical evaluation of the matrix elements are given in the SM [22]. Exchange in the different vortex fusion sectors $\mathds{1}_{\rm is}$ and $\psi$ can be distinguished by computing the exchange phase with doubly-anti-periodic and doubly-periodic fermion boundary conditions, respectively [21]. Indeed, in the anti-periodic case, the even-parity ground state is the quasiparticle vacuum, such that the pair of vortices is created and annihilates in the $\mathds{1}_{\rm is}$ sector. For the periodic case, the FBdG ground state has odd parity and thus the physical even-parity ground state is the first-excited state containing one FBdG fermionic quasiparticle. The vortex pair thus fuses to a fermion. The counterclockwise exchange phases obtained are shown in Fig 3(b), for varying driving amplitude and for different system size. The results converge rapidly with the size towards the prediction for Ising topological order, $R_{\mathds{1}_{\rm is}}^{ee}=e^{-i\pi{C}/8}$ and $R_{\psi}^{ee}=e^{3i\pi C/8}$ for fermion Chern number $C=1$. The $\pi/2$ difference between the exchange phase in the two different sectors supports non-Abelian behaviour [47]. Inverting the sign of the driving phase-shift, $\phi_{y}-\phi_{x}=-\pi/2$, hence inverting the handedness of the circular drive, inverts the sign of the exchange phases. We have thus shown that the zero modes developed by the driven toric code in the high-frequency regime behave as Ising anyons.
@@
Conclusion. With the example of the toric-code model, we have shown that time-periodic driving of an Abelian-anyon system can lead to the emergence of non-Abelian topological order. The ideas presented here may be extended to other 2D topologically ordered systems, building on the property that anyon quasiparticles, being spatially localized, can generically be driven by means of local control, as proposed here. Moreover, they indicate a potential path towards non-Abelian physics, by exploiting out-of-equilibrium Abelian phases as recently realized in engineered quantum systems [6, 7, 5].
@@ Acknowledgements.
The author is grateful to André Eckardt, Nathan Harshman, Alexander Schnell and Isaac Tesfaye for helpful and inspiring discussions.
@@ I Quasiparticle mapping
We use the notation $p(v)$ to indicate the plaquette situated to the immediate top-right of a vertex $v$, and, conversely, $v(p)$ to denote the vertex to the bottom-left of the plaquette $p$. The set of commuting plaquette and vertex operators $\hat{B}_{p}$ and $\hat{A}_{v}$ of the toric code, whose ±$\pm 1$ eigenvalues are associated with the presence or the absence of $e$ and $m$ quasiparticles, represents a complete set of observables on the system Hilbert space on a infinite lattice and an undercomplete one on the torus (since they are not all linearly independent, due to $\prod_{p}\hat{B}_{p}=\prod_{v}\hat{A}_{v}=\mathds{1}$). An equivalent commuting set, sharing the same common eigenstates as $\hat{A}_{v}$ and $\hat{B}_{p}$, is given by the operators $\hat{A}_{v}\hat{B}_{p(v)}$ and $\hat{B}_{p}$, whose ±$\pm 1$ eigenvalues can be associated with the presence of $\psi$ and $e$ quasiparticles instead. A fermion $\psi$ is then conventionally interpreted as located on the plaquette $p$ and composed of a pair of $m$ and $e$ anyons occupying the plaquette $p$ and vertex $v(p)$, respectively (as depicted in Fig. 1). The presence (absence) of a fermion is detected by the $-1$ ($+1$) eigenvalue of $\hat{B}_{p}$, thus counting the parity $(-1)^{n_{p}^{\psi}}$ of the number $n_{p}^{\psi}\in\{0,1\}$ of fermions potentially occupying the plaquette $p$. Similarly, the presence (absence) of an isolated $e$ boson is detected by the $-1$ ($+1$) eigenvalue of $\hat{A}_{v}\hat{B}_{p(v)}$. We thus refer to $\hat{B}_{p}$ as fermion parity operators and to $\hat{A}_{v}\hat{B}_{p(v)}$ as boson parity operators.
The first step of the quasiparticle mapping is to map the toric-code spins to a dual pseudospin space, such that the commuting operators $\hat{A}_{v}\hat{B}_{p(v)}$ and $\hat{B}_{p}$ with ±$\pm 1$ eigenvalues are mapped to commuting pseudospin Pauli-$\hat{Z}$ operators $\hat{\bm{Z}}_{v}^{e}$ and $\hat{\bm{Z}}_{p}^{\varepsilon}$ living on the corresponding vertex and plaquette. The dual pseudospin space is the tensor product $\mathbb{H}_{e}\otimes\mathbb{H}_{\psi}$ of a spin-1/2 space $\mathbb{H}_{e}$ for boson parities and of a spin-1/2 space $\mathbb{H}_{\psi}$ for fermion parities. The pseudospin mapping starts with
$\hat{\bm{Z}}_{v}^{e}=\hat{A}_{v}\hat{B}_{p(v)},\qquad\hat{\bm{Z}}_{p}^{\psi}=\hat{B}_{p}.$ (S6)
On the torus, given the additional constraint that the total $\psi$ and $e$ parities are conserved,
@@
the local parities do not form a complete set. A complete set is obtained by introducing Wilson-loop operators which commute with all local parities, such as
@@
where the paths $\tilde{\Gamma}_{r}$, with $r=x,y$, are depicted in Fig. S1 [when convenient, we use the notation for spin operators defined after Eq. (2) and in Fig. 1(a)]. The Wilson-loop operators describe $\psi$-fermion tunnelling all across the torus in horizontal and vertical direction. Their eigenvalues ±$\pm 1$ distinguish four superselection sectors and, at the end of the mapping, will correspond to the choice of either periodic ($W_{x}=1$ or $W_{y}=1$) or anti-periodic ($W_{x}=-1$ or $W_{y}=-1$) boundary conditions for the fermions. The operators $\hat{W}_{x}$ are mapped to two pseudospins $\hat{\bm{Z}}_{r}^{W}$ in the two-spin space $\mathbb{H}_{W}$, such that the total pseudospin Hilbert space on the torus has the structure $\mathbb{H}_{e}\otimes\mathbb{H}_{\psi}\otimes\mathbb{H}_{W}$. Due to the total fermion and boson parity conservation, single-pseudospin Pauli-$\hat{X}$-type operators which anti-commute with a single $\hat{\bm{Z}}_{v}^{e}$ or $\hat{\bm{Z}}_{p}^{\psi}$ are not physical and cannot be defined: physical operators may only create and annihilate anyon pairs (in an open-boundary system, they can be defined instead by creating a pair with one anyon at the edge [21]). Instead, one can define two-body pseudospin operators, where one of the two pseudospins is conventionally chosen to be located on the first vertex or plaquette on the top-left of the lattice (Fig. S1). These can be defined, for instance, as
$\hat{\bm{X}}_{1}^{e}\hat{\bm{X}}_{v}^{e}=\prod_{i\in\Gamma_{1,v}}\hat{Z}_{i},$ (S9a)
$\hat{\bm{X}}_{1}^{\psi}\hat{\bm{X}}_{p}^{\psi}=\prod_{i\in\Gamma_{1,v}}\hat{Z}_{i}\prod_{j\in\tilde{\Gamma}_{1,p}}\hat{X}_{i},$ (S9b)
@@
where $\Gamma_{1,v}$ is a path connecting vertices 1 and $v$, while $\tilde{\Gamma}_{1,p}$ connects the $1$st and the $p$th plaquette, with the specific path convention depicted in Fig. S1: $\Gamma_{1,v}$ descends vertically from the top-left vertex until the vertex row to which $v$ belongs and the continues horizontally until $v$; $\tilde{\Gamma}_{1,p}$ connects the top-left plaquette to the right-most one in the same row, descends vertically until the plaquette row to which $p$ belongs, and then horizontally to reach $p$. With this choice, it can be verified that spin anti-commutation rules are satisfied by the pseudospins, $\{\hat{\bm{X}}_{1}^{e}\hat{\bm{X}}_{v}^{e},\hat{\bm{Z}}_{v}^{e}\}=0$ and $\{\hat{\bm{X}}_{1}^{\psi}\hat{\bm{X}}_{p}^{\psi},\hat{\bm{Z}}_{p}^{\psi}\}=0$.
Figure S1: Paths used in the quasiparticle mapping of Ref. [21] for the pseudospin mapping ($\Gamma_{x/y}$, $\tilde{\Gamma}_{x/y}$, $\Gamma_{1,v}$ and $\tilde{\Gamma}_{1,p}$) and in the Jordan-Wigner (JW) mapping to fermions.
@@
Using Eqs. (S6) and (I), the two-spin driving terms of the form $\hat{X}_{i}\hat{Z}_{j}$ entering the drive Hamiltonian $\hat{H}_{d}(t)$ of Eq. (2) are mapped to pseudospin operators. The precise mapping depends on whether a term involves spins located on the boundaries, which are intersected by $\Gamma_{r}$ or $\tilde{\Gamma}_{r}$, or not [21]. We report, as an example, only the mapping for terms in the bulk, not intersecting the boundary, which reads
$\hat{X}_{p,R}\hat{Z}_{p,B}=\hat{\bm{X}}_{p}^{\psi}\hat{\bm{X}}_{p+x}^{\psi}$ (S10a)
$\hat{X}_{p,L}\hat{Z}_{p,T}=\left(\prod_{v\in{\mathrm{R}[v(p)]}}\hat{\bm{Z}}_{v}^{e}\right)\left(\prod_{\begin{subarray}{c}p^{\prime}\in\mathrm{L}(p+y)\\ p^{\prime}\in\mathrm{R}(p)\end{subarray}}\hat{\bm{Z}}_{p^{\prime}}^{\psi}\right)\hat{\bm{X}}_{p}^{\psi}\hat{\bm{X}}_{p+y}^{\psi},$ (S10d)
@@
where ${\rm R}(v)$ and $\mathrm{L}(v)$ denote all vertices to the right and to the left of $v$, respectively, and similarly for plaquettes.
@@
$\hat{\bm{Z}}_{v}^{e}=(-1)^{\hat{b}_{v}^{\dagger}\hat{b}_{v}},\quad\hat{\bm{X}}_{v}^{e}=\hat{b}_{v}^{\dagger}+\hat{b}_{v},$ (S11)

### 28 Feb 2024 v2

@@ Abstract
Floquet-Majorana modes. We verify that this Floquet modulation indeed induces topological properties, mimicking a $p+ip$ superconductor, and yields Floquet-Majorana modes. We study these properties by numerically computing the Floquet Bogoliubov-de Gennes (BdG) Hamiltonian $\hat{H}_{F,\mathrm{BdG}}$, defined by $\hat{H}_{F}=(1/2)(\hat{\bm{f}}^{\dagger},\hat{\bm{f}})\hat{H}_{F,\mathrm{BdG}}(\hat{\bm{f}},\hat{\bm{f}}^{\dagger})^{T}$, as well as its quasienergy spectrum and ground state parity, from the driven model of Eq. (3[25]. The topological phase is expected for $|\mu|<4J$ and is characterized by a non-zero Chern number of the negative-energy fermion band [16, 27, 26]. Tuning the frequency $\omega$ to this topological region ($\mu\simeq-2J$), in the absence of $e$ vortices [$\mathcal{S}_{p}^{(r)}=1$], we find four nearly-degenerate high-frequency ‘ground’ states on a torus, one in each superselection sector corresponding to periodic or anti-periodic boundary conditions in $x$ and $y$ direction [16]. The topological degeneracy is approached exponentially with increasing system size, as shown in Fig. (2)(a), and only three ground states have even fermion parity and are thus physical, signalling three Ising-anyon types $e,\mathds{1}_{\rm is},\psi$. On a cylinder, the quasienergy spectrum exhibits two edge modes [Fig. 2(b)], indicating the development of nontrivial topology and a non-zero Chern number.
@@
Introducing a pair of separated $e$ particles on the torus, approximate Majorana modes at near-zero quasienergy appear [Fig. 2(c)], which disappear when $\omega$ is detuned out of the topological phase. The high-frequency description of Eq. (5) thus correctly captures the topological properties of the system, which are not spoiled by the higher-order contributions to the high-frequency expansion neglected in (5). Increasing the driving amplitudes $J=\Delta$, while adapting $\mu=-2J$ to stay in the topological region, the fermions enter an anomalous Floquet phase [36, 37, 30]. Although the time-averaged description of Eq. (5) is no longer valid as they depart from the high-frequency limit, the system still exhibits Majorana modes, at both quasienergy zero and ±$\pm\pi/T$, see Fig. 2(c). This effect, only possible in driven systems [38, 39, 36, 40, 41, 42, 43, 44, 45], has been proposed as a means to realize non-Abelian braiding and quantum computation in 1D Kitaev quantum wires [46, 47, 48, 49].
@@
Non-Abelian exchange phases. We analyze whether the exchange statistics of two $e$ particles (vortices) in the fermion topological phase is non-Abelian, as predicted in the presence of Majorana modes [8, 28]. Two vortices are exchanged explicitly in different fusion sectors following Levin-Wen’s protocol [50, 8, 51], sketched in Fig. 2(d), and Ref. [24]. The protocol computes the difference in the Berry phase accumulated by moving vortices along two paths $P$ and $P^{\prime}$, sharing the same set of positions but involving a different order of vortex moves. This guarantees that the resulting phase is only determined by the exchange statistics [50]. For each intermediate vortex configuration |$\ket{\vec{n}_{i}^{e}}$ along the paths, we numerically determine the corresponding even-parity ground state |$\ket{\Phi_{i}}$ of the Floquet-BdG fermion Hamiltonian [25]. The Berry phase accumulated along the path $P$ is computed from the sequence of ground states as $\theta_{P}=\mathrm{Arg}\prod_{(i,i+1)\in P}\bra{\Phi_{i+1}}\hat{\tilde{S}}^{(r)}_{i+1,i}\ket{\Phi_{i}}$. Here, $\hat{\tilde{S}}^{(r)}_{i+1,i}$ is the fermionic part of the quasiparticle representation of the spin operator $\hat{Z}_{i}$ which converts the vortex configuration |$\ket{\vec{n}^{e}_{i}}$ into |$\ket{\vec{n}^{e}_{i+1}}$ by displacing one $e$ particle. Indeed, the Pauli operators $\hat{Z}_{v,R}$ and $\hat{Z}_{v,B}$ between vertices $v$ and $v+r$ (with $r=x$ or $y$), map to operators $\hat{t}_{v}^{(r)}=\hat{\tilde{\mathcal{S}}}_{v}^{(r)}(\hat{b}_{v}^{\dagger}+\hat{b}_{v})(\hat{b}_{v+r}^{\dagger}+\hat{b}_{v+r})$, describing boson motion but also containing a fermionic part to capture semion statistics: $\hat{\tilde{\mathcal{S}}}_{v}^{(x)}=1$ and $\hat{\tilde{\mathcal{S}}}_{v}^{(y)}=\prod_{p\in\mathrm{L}[p(v+y)]}(-1)^{\hat{f}_{p}^{\dagger}\hat{f}_{p}}$, where $\mathrm{L}[p(v+y)]$ denotes all plaquettes to the left of the spin between $v$ and $v+y$. Details about the numerical evaluation of the matrix elements in $\theta_{P}$ are given in the SM [25]. We compare the exchange phase in different fusion sectors by creating the $e$-particle pair on top of two different degenerate ground states. The prediction for Ising anyons is that the exchange phase $R_{\mathds{1}_{\rm is}}^{\sigma\sigma}$ in the $\mathds{1}_{\rm is}$ sector (where the vortices fuse to the vacuum) and $R_{\psi}^{\sigma\sigma}$ in the $\psi$ sector (where they fuse to a fermion $\psi$) differ by $\pi C/2$ for Chern number $C$, namely $R_{\mathds{1}_{\rm is}}^{\sigma\sigma}=e^{-i\pi{C}/8}$ and $R_{\psi}^{\sigma\sigma}=e^{3i\pi C/8}$ [8]— a signature of their non-Abelian nature [52]. The two sectors $\mathds{1}_{\rm is}$ and $\psi$ are selected by considering doubly-antiperiodic and doubly-periodic fermion boundary conditions, respectively [24]. In the antiperiodic case, the even-parity ground state is the quasiparticle vacuum, such that the pair of vortices is created from and fuses to the vacuum. For the periodic case, the BdG ground state has odd parity and thus the physical even-parity ground state is the first-excited state containing one BdG fermion. The vortex pair thus fuses to this fermion. The counterclockwise exchange phases obtained are shown in Fig 2(e), for varying driving amplitude in the high-frequency regime and for different system size. The results converge rapidly with the size towards the prediction for Ising topological order with $C=1$. Inverting the handedness of the circular drive sign, $\phi_{y}-\phi_{x}=-\pi/2$, inverts the sign of the exchange phases. In conclusion, we have shown that the $e$ vortices carrying Floquet-Majorana modes, arising in the driven toric code in the high-frequency regime, behave like non-Abelian Ising anyons.
@@
Conclusion. With the example of the toric-code model, we have shown that time-periodic driving of an Abelian-anyon system can lead to the emergence of non-Abelian topological order. The ideas presented here may be extended to other 2D topologically-ordered systems, exploiting the fact that pointlike anyon quasiparticles can be driven through local control. Moreover, they indicate a potential path towards non-Abelian physics, building on out-of-equilibrium Abelian phases as recently realized in engineered quantum systems [6, 7, 5].
@@ Acknowledgements.
The author is grateful to André Eckardt, Nathan Harshman, Alexander Schnell, Isaac Tesfaye and Sandro Wimberger for helpful and inspiring discussions. The author also acknowledges funding by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) via the Research Unit FOR 2414 - project number No. 277974659.
@@ I Quasiparticle mapping
We use the notation $p(v)$ to indicate the plaquette situated to the immediate top-right of a vertex $v$, and, conversely, $v(p)$ to denote the vertex to the bottom-left of the plaquette $p$. The set of commuting plaquette and vertex operators $\hat{B}_{p}$ and $\hat{A}_{v}$ of the toric code, whose ±$\pm 1$ eigenvalues are associated with the presence or the absence of $e$ and $m$ quasiparticles, represents a complete set of observables on the system Hilbert space on a infinite lattice. For an $L\times L$ lattice on a torus, hosting $2L^{2}$ spins and $L^{2}$ vertices and plaquettes, they are under-complete because of the condition $\prod_{p}\hat{B}_{p}=\prod_{v}\hat{A}_{v}=\mathds{1}$: Specifying all their (independent) plaquette and vertex eigenvalues imposes $2^{(L^{2}-1)}\cdot 2^{(L^{2}-1)}$ constraints in the $2^{L^{2}}$ Hilbert, leaving two spin-1/2 degrees of freedom free corresponding to the topological degeneracy. An equivalent commuting set, sharing the same common eigenstates of the operators $\hat{A}_{v}$ and $\hat{B}_{p}$, is given by the operators $\hat{A}_{v}\hat{B}_{p(v)}$ and $\hat{B}_{p}$, whose ±$\pm 1$ eigenvalues can be associated with the presence of $e$ and $\psi$ quasiparticles instead. A fermion $\psi$ is then conventionally interpreted as located on the plaquette $p$ and composed of a pair of $m$ and $e$ anyons occupying the plaquette $p$ and vertex $v(p)$, respectively (as depicted in Fig. 1). The presence (absence) of a fermion is detected by the $-1$ ($+1$) eigenvalue of $\hat{B}_{p}$, thus counting the parity $(-1)^{n_{p}^{\psi}}$ of the number of fermions $n_{p}^{\psi}\in\{0,1\}$ potentially occupying the plaquette $p$. Similarly, the presence (absence) of an isolated $e$ boson is detected by the $-1$ ($+1$) eigenvalue of $\hat{A}_{v}\hat{B}_{p(v)}$. We thus refer to $\hat{B}_{p}$ as fermion parity operators and to $\hat{A}_{v}\hat{B}_{p(v)}$ as boson parity operators.
Figure S1: Paths used in the quasiparticle mapping of Ref. [24] (a) for the pseudospin mapping ($\Gamma_{x/y}$, $\tilde{\Gamma}_{x/y}$, $\Gamma_{1,v}$ and $\tilde{\Gamma}_{1,p}$) and (b) in the Jordan-Wigner (JW) mapping to fermions.
The first step of the quasiparticle mapping is to map the toric-code spins to a dual pseudospin space, such that the commuting operators $\hat{A}_{v}\hat{B}_{p(v)}$ and $\hat{B}_{p}$ with ±$\pm 1$ eigenvalues are mapped to commuting spin-1/2 Pauli-$\hat{Z}$ operators $\hat{\bm{Z}}_{v}^{e}$ and $\hat{\bm{Z}}_{p}^{\psi}$ living on the corresponding vertex and plaquette. The dual pseudospin space is the tensor product $\mathbb{H}_{e}\otimes\mathbb{H}_{\psi}$ of a space $\mathbb{H}_{e}$ for boson parities and of a space $\mathbb{H}_{\psi}$ for fermion parities. The pseudospin mapping starts with
$\hat{A}_{v}\hat{B}_{p(v)}\leftrightarrow\hat{\bm{Z}}_{v}^{e},\qquad\hat{B}_{p}\leftrightarrow\hat{\bm{Z}}_{p}^{\psi},$ (S6)
such that the toric-code Hamiltonian is mapped to
$\hat{H}_{\rm tc}\leftrightarrow-\frac{g}{2}\sum_{v}\hat{\bm{Z}}_{v}^{e}\hat{\bm{Z}}_{p(v)}^{\psi}-\frac{g}{2}\sum_{p}\hat{\bm{Z}}_{p}^{\psi}\ .$ (S7)
On the torus, given the additional constraint
@@
the total $\psi$ and $e$ parities are conserved and the local parities do not form a complete set of operators. A complete set is obtained by introducing Wilson-loop operators which commute with all local parities, such as
@@
where the paths $\tilde{\Gamma}_{r}$, with $r=x,y$, are depicted in Fig. S1 and the operators $\hat{W}_{x}$ and $\hat{W}_{y}$ are depicted in Fig. S2(b) [when convenient, we use the notation for spin operators defined after Eq. (2) and in Fig. 1(a)]. The Wilson-loop operators describe $\psi$-fermion tunnelling all across the torus in horizontal and vertical direction. Their eigenvalues ±$\pm 1$ distinguish four superselection sectors and, at the end of the mapping, will correspond to the choice of either periodic ($W_{x}=1$ or $W_{y}=1$) or anti-periodic ($W_{x}=-1$ or $W_{y}=-1$) boundary conditions for the fermions. The operators $\hat{W}_{x}$ are mapped to two pseudospins $\hat{\bm{Z}}_{r}^{W}$ in the two-spin space $\mathbb{H}_{W}$, such that the total pseudospin Hilbert space on the torus has the structure $\mathbb{H}_{e}\otimes\mathbb{H}_{\psi}\otimes\mathbb{H}_{W}$. Due to the total fermion and boson parity conservation, single-pseudospin Pauli-$\hat{X}$-type operators which anti-commute with a single $\hat{\bm{Z}}_{v}^{e}$ or $\hat{\bm{Z}}_{p}^{\psi}$ are not physical and have no mapping in terms of operators on the original spin system: physical spin operators may only create and annihilate anyon pairs, since they always anticommute with an even number of local parities. However, one can define two-body pseudospin operators, anticommuting with $\hat{\bm{Z}}_{v}^{e}$ or $\hat{\bm{Z}}_{p}^{\psi}$, where one of the two pseudospins is conventionally chosen to be located on the first vertex or plaquette on the top-left of the lattice, which we label as $v=1$ and $p=1$, respectively [Fig. S1(a) and Fig. S2(a)]. These can be defined, for instance, as
$\hat{\bm{X}}_{1}^{e}\hat{\bm{X}}_{v}^{e}\leftrightarrow\prod_{i\in\Gamma_{1,v}}\hat{Z}_{i},$ (S10a)
$\hat{\bm{X}}_{1}^{\psi}\hat{\bm{X}}_{p}^{\psi}\leftrightarrow\prod_{i\in\Gamma_{1,v}}\hat{Z}_{i}\prod_{j\in\tilde{\Gamma}_{1,p}}\hat{X}_{j},$ (S10b)
@@
where $\Gamma_{1,v}$ is a path connecting vertices 1 and $v$, while $\tilde{\Gamma}_{1,p}$ connects the $1$st and the $p$th plaquette, with the specific path convention depicted in Fig. S1(a): $\Gamma_{1,v}$ descends vertically from the top-left vertex until the vertex row to which $v$ belongs and then continues horizontally until $v$; $\tilde{\Gamma}_{1,p}$ connects the top-left plaquette to the right-most one in the same row, descends vertically until the plaquette row to which $p$ belongs, and then horizontally to reach $p$. With this choice, it can be verified that spin anti-commutation rules are satisfied by the pseudospins, $\{\hat{\bm{X}}_{1}^{e}\hat{\bm{X}}_{v}^{e},\hat{\bm{Z}}_{v}^{e}\}=0$ and $\{\hat{\bm{X}}_{1}^{\psi}\hat{\bm{X}}_{p}^{\psi},\hat{\bm{Z}}_{p}^{\psi}\}=0$. The operators in Eq. (I) are depicted in Fig. S2(a).
@@
Using Eqs. (S6) and (I), the two-spin driving terms of the form $\hat{X}_{i}\hat{Z}_{j}$ entering the drive Hamiltonian $\hat{H}_{d}(t)$ of Eq. (2) are mapped to pseudospin operators. The precise mapping depends on whether a term involves spins located on the boundaries, which are intersected by $\Gamma_{r}$ or $\tilde{\Gamma}_{r}$, or not [24]. We only report, as an example, the mapping for terms in the bulk, not intersecting the boundary, which reads
$\hat{X}_{p,R}\hat{Z}_{p,B}\leftrightarrow$ $\hat{\bm{X}}_{p}^{\psi}\hat{\bm{X}}_{p+x}^{\psi}$ (S11a)
$\hat{X}_{v(p),R}\hat{Z}_{v(p),B}\leftrightarrow$ $\Big{[}\prod_{v^{\prime}\in{\mathrm{R}[v(p)]}}\hat{\bm{Z}}_{v^{\prime}}^{e}\Big{]}$
×$\times\Big{[}\prod_{\begin{subarray}{c}p^{\prime}\in\mathrm{L}(p+y)\\ p^{\prime}\in\mathrm{R}(p)\end{subarray}}\hat{\bm{Z}}_{p^{\prime}}^{\psi}\Big{]}\hat{\bm{X}}_{p}^{\psi}\hat{\bm{X}}_{p+y}^{\psi},$ (S11d)
@@
where ${\rm R}(v)$ and $\mathrm{L}(v)$ denote all vertices to the right and to the left of $v$, respectively, and similarly for plaquettes. The mapping of Eq. (S11a) for horizontal processes is sketched in Fig. S2(c), while the mapping of Eq. (S11d) for vertical processes is sketched in Fig. S3(a)-(b).
Figure S2: (a) Strings of spin operators mapping to the two-body pseudospin operators $\hat{\bm{X}}_{1}^{\psi}\hat{\bm{X}}_{p}^{\psi}$ and $\hat{\bm{X}}_{1}^{e}\hat{\bm{X}}_{v}^{e}$. (b) Strings of spin operators realizing the Wilson loop operators of Eq. (S9). (c) Sketch of the mapping from the two-spin terms $\hat{Z}_{p,B}\hat{X}_{p,R}$, inducing horizontal tunnelling and pairing of fermions on plaquettes $p$ and $p+x$, to two-body pseudospin operators $\hat{\bm{X}}_{p}^{\psi}\hat{\bm{X}}_{p+x}^{\psi}$.
@@
$\hat{\bm{Z}}_{v}^{e}\leftrightarrow(-1)^{\hat{b}_{v}^{\dagger}\hat{b}_{v}},\quad\hat{\bm{X}}_{v}^{e}\leftrightarrow\hat{b}_{v}^{\dagger}+\hat{b}_{v},$ (S12)